OpenStudy (anonymous):

-46 times a number minus 46 is equal to 77 less than a number. what is the number

OpenStudy (vishal_kothari):

is it a rational number?

OpenStudy (anonymous):

-46*x-46=77-x x = -41/15

OpenStudy (anonymous):

-46x-46=x-77 -47x=-31 x=31/47

OpenStudy (anonymous):

I'm wrong

OpenStudy (anonymous):

whats the equation

OpenStudy (anonymous):

-46x-46=x-77

OpenStudy (vishal_kothari):

yeah...